A cylinder has a diameter of 12 in. and a height of 6 in. Find the volume.
Answer by TimothyLamb(4379)

v = pi * rr * h
radius = diameter/2
v = pi * dd/4 * h
---
v = 3.1415 * 12*12/4 * 6
v = 678.564 cubic inches
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php